One-third of a number is 2 more than one-sixth of the same number. What is the number?
Answer by checkley77(12569)

x/3=x/6+2
x/3-x/6=2
(2x-x)/6=2
x=2*6
x=12 ans.Proof:
12/3=12/6+2
4=2+2
4=4